OpenStudy (anonymous):

(3w-6/9) = (3) - (w-3/3)

OpenStudy (anonymous):

first reduce your fractions so that 6/9 = 2/3 and 3/3 = 1 3w - 2/3 = 3 - (w-1) now think of the minus sign after your 3 as a -1 and multiply that by both the w and the -1 in your parentheses. 3w -2/3 = 3 -w +1 add your 3 and 1 3w -2/3 = 4 - w add w to both sides 4w - 2/3 = 4 next you'll need to add 2/3 to each side. To do this you'll need a common denominator of 3. Rewrite 4 as 12/3. 4w=12/3 + 2/3 add 4w = 14/3 Now you need to divide by 4 so that your w is alone on that side of the equation. A trick is to remember that dividing by a number is the same as multiplying by the reciprocal. That means that\[4w \div 4\] is the same as \[4w \times 1/4\]. So if you multiply both sides by 1/4 the 4s will cancel out by the w and then to multiply 14/3 and 1/4 you just multiply straight across to get 14/12. w=14/12 which can be reduced to either the improper 7/6 or the mixed fraction 1 and 1/6

OpenStudy (anonymous):

 is this how you had it set up?